OpenStudy (anonymous):

Jake is comparing the prices of two mattress cleaning companies. Company A charges $30 per mattress and an additional $13 as service charges. Company B charges $28 per mattress and an additional $15 as service charges. Part A: Write equations to represent Company A and Company B's total mattress cleaning charges for a certain number of mattresses. Define the variable used in the equations. Part B: Which company would charge less for cleaning 4 mattresses? Justify your answer. Part C: How much money is saved by using the services of Company B instead of Company A to cle

OpenStudy (anonymous):

part c ;Part C: How much money is saved by using the services of Company B instead of Company A to clean 7 mattresses?

OpenStudy (ranga):

Let N be the number of mattresses to be cleaned and C be the total cost. For company A: C = 30N + 13 For company B: C = 28N + 15

OpenStudy (ranga):

For cleaning 4 mattresses, put N = 4 Company A: C = 30 * 4 + 13 = $133 Company B: C = 28 * 4 + 15 = $127 Therefore, for cleaning 4 mattresses, Company B will charge less.

OpenStudy (ranga):

For part C, put N = 7 and calculate the total cost with Company A and Company B and find the difference.

OpenStudy (anonymous):

ty :)

OpenStudy (ranga):

yw. For c) I I am getting $12 as the money saved in using Company B to clean 7 mattresses.